# Understanding Bias Reinforcement in LLM Agents Debate

### v4

@@ 7 Conclusions
Impact Statement
This work advances Machine Learning by enhancing LLMs’ strategic reasoning through our approach, addressing bias reinforcement and lack of perspective diversity in Multi-Agent Debate (MAD). Our method improves decision-making in adversarial settings, with potential applications in automated negotiations, economics, and multi-agent systems. However, stronger AI-driven strategies could be misused in manipulative or deceptive contexts. To ensure ethical deployment, future research should focus on integrating fairness constraints and transparency mechanisms in AI decision-making.
@@ Appendix A MetaNIM Arena
A.1 Theoretical Background on Combinatorial Impartial Games
Each game in MetaNIM Arena is a combinatorial impartial game. We begin by outlining the relevant theoretical foundation, starting with the Sprague–Grundy theorem.
A.1.1 Sprague-Grundy Theorem
The Sprague-Grundy theorem provides a fundamental method for analyzing impartial combinatorial games by decomposing complex games into simpler, independent subgames. As discussed in Section 3.4, every impartial combinatorial game can be represented as a directed acyclic graph (DAG). However, directly computing Grundy numbers recursively from terminal states is often impractical.
@@ A.1.1 Sprague-Grundy Theorem
Let $\mathcal{G}_{1}=(X_{1},F_{1}),\mathcal{G}_{2}=(X_{2},F_{2}),\dots,\mathcal{G}_{n}=(X_{n},F_{n})$ be DAGs representing $n$ impartial combinatorial games. The disjunctive sum of $\mathcal{G}_{1},\dots,\mathcal{G}_{n}$ is a DAG $\mathcal{G}=(X,F)$ defined as follows:
@@
This result implies that Grundy numbers for complex games, such as Kayles or Chomp, can be efficiently computed by decomposing them into simpler subgames and combining the Grundy numbers using bitwise XOR.
For example, consider a variant of the game Kayles played on two separate rows of pins, each forming an independent subgame. Suppose we computed the Grundy numbers separately for these rows, obtaining Grundy numbers $7$ for the first row and $4$ for the second row. By the Sprague-Grundy theorem, the combined Grundy number of the position is given by:
@@
A.1.2 Basic Discussions on the Optimal Strategy
@@
[…]
Concretely, if $G(S)=0$, then no valid move leads to a successor $S^{\prime}$ with $G(S^{\prime})=0$. In other words, from $S$, the player to move cannot transition the game into a $G(\cdot)=0$ state. Because a state $G(S^{\prime})=0$ corresponds to a losing position for the player who faces it, the mover in state $S$ has no way to force the opponent into a losing position on the next turn. Hence, $S$ is losing for the player to move.
@@
A.2 Game Variants in MetaNIM Arena
A.2.1 NIM
Nim is a mathematical strategy game where two players alternate turns removing objects from distinct heaps/piles. The classic version follows these rules:
  • Heaps: The game starts with $k$ heaps containing $n_{1},n_{2},\dotsc,n_{k}$ objects respectively
  • Moves: On their turn, a player must remove at least 1 to previously fixed number of objects from exactly one heap
  • Objective: The player who takes the last remaining object wins (normal play convention)
Mathematical Strategy
The game can be analyzed using binary representations through the concept of Nimbers (Grundy numbers). For any position, the key is to calculate the binary XOR (exclusive OR) sum of all heap sizes:
$\text{Nim-sum}=n_{1}\oplus n_{2}\oplus\dotsb\oplus n_{k}$
A position is losing if the Nim-sum equals 0. The winning strategy consists of always moving to a position with Nim-sum 0. For the single-heap variant (as in our current game), this simplifies to maintaining modular arithmetic conditions.
Example
Consider a game with heaps [3, 4, 5]:
$3$ $=011_{2}$
$4$ $=100_{2}$
$5$ $=101_{2}$
Nim-sum $=011_{2}\oplus 100_{2}\oplus 101_{2}=010_{2}=2\neq 0$
The first player can win by removing 2 objects from the 5-object heap to make the new Nim-sum 0.
Variants
Several Nim variants exist, including:
  • Single-heap Nim (as in our current game)
  • Misère Nim (player taking last object loses)
  • Multi-heap Nim with different removal constraints
The fundamental mathematical principles of combinatorial game theory apply to all variants.
A.2.2 Fibonacci
The Fibonacci Game, also known as Fibonacci Nim, is a combinatorial number game where players alternate removing items from a pile, with move constraints based on the Fibonacci sequence. The rules are:
  • Initial Move: First player takes $1\leq k<n$ items from a pile of $n$ items
  • Subsequent Moves: Each player must take between 1 and twice the number of items taken by their opponent in the previous move
  • Objective: The player who takes the last item wins
Mathematical Strategy
The game is governed by Fibonacci numbers ($F_{1}=1,F_{2}=2,F_{n}=F_{n-1}+F_{n-2}$) and Zeckendorf’s Theorem, which states that every positive integer can be uniquely expressed as a sum of non-consecutive Fibonacci numbers.
  • Losing Positions: Pile sizes equal to Fibonacci numbers ($F_{n}$)
  • Winning Strategy: Reduce the pile to the largest Fibonacci number smaller than the current size
For a pile of size $m$, its Zeckendorf representation is:
$m=F_{k_{1}}+F_{k_{2}}+\dots+F_{k_{r}}\quad(|k_{i}-k_{j}|\geq 2)$
The optimal first move is to remove the smallest Fibonacci number in this decomposition.
Example
For a starting pile of $m=20$:
$\text{Zeckendorf: }20$ $=13+5+2\quad(F_{7}=13,F_{5}=5,F_{3}=2)$
First move $=\text{Remove smallest term }2$
New pile $=18=13+5$
Now the opponent faces a position composed purely of Fibonacci numbers. Any move they make ($1\leq x\leq 4$) can be countered by reducing the pile to the next Fibonacci number.
Key Properties
  • If $m$ is a Fibonacci number, the first player will lose against perfect play
  • The number of moves in a game is always $\leq$ the index of the largest Fibonacci number $\leq m$
  • The golden ratio $\phi=\frac{1+\sqrt{5}}{2}$ emerges in win/loss probability analysis
Variants
  • Reverse Fibonacci Nim (last player to move loses)
  • Multi-pile Fibonacci games
  • Constrained Fibonacci sequences (e.g., Tribonacci variants)
This game demonstrates deep connections between combinatorial game theory, number theory, and the Fibonacci sequence.
A.2.3 Kayles
Kayles is an impartial combinatorial game played with a linear arrangement of pins where players alternate knocking down pins under specific adjacency rules. First analyzed in 1929 by Dudeney and later studied by Conway and Berlekamp, it demonstrates complex mathematical patterns.
Basic Rules
  • Initial Setup: A row of $n$ identical pins
  • Moves: On each turn, a player must either:
    • Knock down 1 pin
    • Knock down 2 adjacent pins
  • Objective: Last player to make a valid move wins (normal play convention)
Mathematical Analysis
The game is analyzed using Grundy numbers and the Sprague-Grundy theorem. Positions split into independent segments after moves create disjunctive game components.
  • Let $G(n)$ be the Grundy number for a row of $n$ pins
  • Recursive Grundy number calculation:
    $G(n)=\mathrm{mex}\{G(n-1),G(n-2),G(a)\oplus G(b)\}$
    where a+b=nka+b=n-k for $k\in\{1,2\}$, and $\mathrm{mex}$ = minimum excludant
Key Patterns
  • Positions with Grundy number 0 are losing positions
  • The Grundy sequence becomes periodic with period 12 for large $n$
  • Known solution: $G(n)=n\mod 12$ when $n\geq 70$
Example
Consider a row of 4 pins:
$G(0)$ $=0$
$G(1)$ $=\mathrm{mex}\{G(0)\}=1$
$G(2)$ $=\mathrm{mex}\{G(1),G(0),G(0)\oplus G(0)\}=\mathrm{mex}\{1,0,0\}=2$
$G(3)$ $=\mathrm{mex}\{G(2),G(1),G(1)\oplus G(0)\}=\mathrm{mex}\{2,1,1\}=0$
$G(4)$ $=\mathrm{mex}\{G(3),G(2),G(2)\oplus G(0)\}=\mathrm{mex}\{0,2,2\}=1$
A row of 4 pins has Grundy number 1, making it a winning position.
Strategic Principles
  • Split long rows into independent segments with XOR-sum 0
  • Mirror opponent’s moves in symmetric positions
  • Avoid leaving isolated single pins
Variants
  • Circular Kayles (pins arranged in a circle)
  • Multi-row Kayles
  • $k$-Kayles (allow knocking down up to $k$ adjacent pins)
  • Misère Kayles (last player to move loses)
Computational Complexity
Kayles is:
  • PSPACE-complete for general positions
  • Solved in linear time for standard single-row play
  • Used in complexity theory to study impartial games
This analysis demonstrates how simple rule sets can generate complex mathematical structures. The complete Grundy number sequence for Kayles was only fully determined through extensive computational analysis.
A.2.4 Chomp
Chomp is an impartial combinatorial game first formulated by David Gale in 1974. Played on a rectangular grid representing a chocolate bar, it features unique topological constraints and demonstrates fundamental principles of partially ordered sets (posets).
Basic Rules
  • Initial Setup: An $m\times n$ rectangular grid of ”chocolate squares”
  • Special Square: The lower-left square (position (1,1)) is poisoned
  • Moves: On each turn, a player must:
    • Select any remaining square
    • Remove (”chomp”) all squares above and/or to the right of the selected square
  • Objective: Avoid taking the poisoned square - last player to make a valid move wins (normal play convention)
Mathematical Analysis
Chomp is particularly significant in combinatorial game theory because:
  • It is a partisan game with inherent asymmetry
  • The starting position is a poset under component-wise ordering
  • A winning strategy exists for the first player (proven by strategy-stealing argument), though explicit strategies are unknown for most grid sizes
Key Theorem (Gale, 1974)
Theorem A.3.
For any initial grid size $m\times n$ where $m,n\geq 2$, the first player has a winning strategy.
Example: 2 × 3 Grid
[…]
First player wins by taking the (2,3) square:
  • If Player 2 takes (1,3), Player 1 takes (2,2)
  • If Player 2 takes (2,2), Player 1 takes (1,2)
  • All paths eventually force Player 2 to take the poison
Strategic Principles
  • Maintain control of the antidiagonal
  • Force symmetry when possible
  • Reduce the game to independent subgames
  • Avoid leaving isolated columns
Computational Complexity
  • General Chomp is PSPACE-complete
  • Solved in polynomial time for:
    • 2 × n grids
    • Square grids up to 5 × 5
  • Number of winning positions grows exponentially with grid size
Variants
  • 3D Chomp (cuboidal grids)
  • Circular Chomp
  • Hypergraph Chomp
  • Misère Chomp (taking poison square wins)
  • Numerical Chomp (played on factor lattices)
Significance
Chomp demonstrates fundamental connections between:
  • Combinatorial game theory
  • Computational complexity
  • Algebraic geometry (via Gröbner basis interpretations)
Despite its simple rules, Chomp remains unsolved for general grid sizes, making it an active research area in computational combinatorics.
A.2.5 Corner Queen
Corner Queen is a two-player combinatorial impartial game played on a rectangular grid, where a single queen starts at an arbitrary position and players alternate moving it toward the bottom-left corner. The player who moves the queen to the target wins.
Basic Rules
  • Initial Setup: A queen is placed on an $m\times n$ grid at position $(x,y)$
  • Moves: On each turn, a player moves the queen in one of the following directions:
    • Left: $(x,y)\rightarrow(x-k,y)$
    • Down: $(x,y)\rightarrow(x,y-k)$
    • Diagonally down-left: $(x,y)\rightarrow(x-k,y-k)$
    • for any $k\in\mathbb{Z}_{>0}$ such that the move stays within the board
  • Objective: The player who moves the queen to position $(0,0)$ wins
Mathematical Analysis
Corner Queen is mathematically equivalent to Wythoff’s Game, a well-studied impartial game in combinatorial game theory. Each game state $(x,y)$ corresponds to a position on the board where $x,y\in\mathbb{N}$ and $x\leq y$ (without loss of generality).
  • The game’s Grundy function $G(x,y)$ satisfies analyzed via Beatty sequences for Wythoff’s Game
  • The losing positions (also called $P$-positions) are given by pairs of the form:
    $(\lfloor k\phi\rfloor,\lfloor k\phi^{2}\rfloor),\quad k\in\mathbb{N}$
    where $\phi=\frac{1+\sqrt{5}}{2}$ is the golden ratio
  • Examples: The first few $P$-positions are $(1,2)$, $(3,5)$, $(4,7)$, $(6,10)$, $(8,13)$, and $(9,15)$.
Optimal Strategy
  • A position is losing if and only if it lies on the Wythoff pairs described above
  • The winning strategy is to move the queen to the nearest $P$-position
  • These positions are sparse and non-periodic, but can be computed efficiently using Beatty sequences
A.3 Our Dataset used in Experiments
We construct simple dataset based on the MetaNIM Arena. This dataset doesn’t require any opponent model because samples in this dataset is focusing on the specific scene in each game.
@@ A.3 Our Dataset used in Experiments
Action space 1–3 dynamic (max 30) $x,y$ coordinate (scenario-dependent) single pin or two adjacent pins
@@
A.4 Our Simulator used in Experiments
We build simulator based on the MetaNIM Arena. This simulator requires any opponent available to receive prompt and make an output as an action. Here, we utilize the gpt-4o model as an opponent.
@@ A.4 Our Simulator used in Experiments
Appendix B Prompts Design
B.1 Game Prompts
Table 10: NIM game basic input prompt.
#Game Role:
You are {agent[’name’]}, a participant in a game of Nim variants.
#Objective:
Your goal is to win the game by taking all remaining items on your turn, leaving no items for your opponent. The person who takes the last item wins.
#Game Rule:
There is a single pile of items. You can take between 1 and {max_take} items on your turn.
#Current State:
There are {remaining_items} items remaining in the pile.
#Task:
Based on the current state of the game, decide how many items you will take (between 1 and {max_take}) on this turn.
Table 11: Fibonacci game basic input prompt.
#Game Role:
You are {agent[’name’]}, a participant in a simple Fibonacci game.
#Objective:
Your goal is to win the game by taking all remaining stones on your turn, leaving no stones for your opponent. The person who takes the last stones wins.
#Game Rule:
1. There is a single pile of stones.
@@ B.1 Game Prompts
#Current State:
There are {remaining_items} stones remaining in the pile.
#Task:
You are the first player. Based on the current state of the game, decide how many items you will take (between 1 and {remaining_items - 1}) on this turn.
Table 12: Kayles game basic input prompt.
#Game Role:
You are {agent[’name’]}, a participant in a game of Kayles.
#Objective:
Your goal is to win the game by leaving your opponent with no valid moves. The player who takes the last pin(s) wins.
#Game Rule:
1. There is a single row of pins.
2. On your turn, you can remove: 1 pin, 2 adjacent pins. 3. You cannot remove non-adjacent pins or pins that have already been removed.
#Current State:
The row of pins is represented as a binary string:

### v5

@@ 8 Conclusion
Acknowledgments
This work was supported by Center for Applied Research in Artificial Intelligence (CARAI) grant funded by Defense Acquisition Program Administration (DAPA) and Agency for Defense Development (ADD) (UD230017TD).
We sincerely thank the action editor, Antoine Bosselut, for their guidance and careful handling of the review process. We also thank the anonymous reviewers for their thoughtful and constructive feedback, which substantially improved the clarity and quality of this work.
@@ Appendix A MetaNIM Arena
A.1 Background on Impartial Games
Games in the MetaNIM Arena are impartial games. We begin by outlining the relevant theoretical foundation, starting with the Sprague–Grundy theorem. In Table 4, we present the overall rules of the Nim game and its optimal strategies.
A.1.1 Theory and Strategy
All MetaNIM Arena games are impartial games, forming a Directed Acyclic Graph (DAG) where vertices represent game states and edges denote valid moves. The Grundy Number framework, along with the Sprague-Grundy Theorem, guarantees the existence of a winning strategy and provides a concrete method to determine it. In the MetaNIM Arena, each state has a mathematically provable optimal move, allowing an LLM’s decisions to be evaluated against the theoretical optimal strategy—a key advantage for unbiased assessment.
Definition A.1 (Grundy Number).
For a finite impartial game under normal play (where the last player to make a valid move wins), the Grundy number (or Nimber) $G(S)$ is recursively defined as follows. If $S$ is a terminal state with no valid moves, set $G(S)=0$. Otherwise,
$G(S)\;=\;\mathrm{mex}\{\,G(S^{\prime})\;|\;S^{\prime}\!\text{ is reachable from }S\,\}\,.$
Here, $\mathrm{mex}(X)$ is the smallest nonnegative integer not in $X$.
Note that the Grundy number is well-defined for every impartial game, since the game’s state space forms a DAG. In many impartial games, direct enumeration of all possible move sequences is computationally infeasible. However, with Sprague-Grundy Theorem, we can easily calculate Grundy Numbers on complex games. See Appendix A.1.2 for further discussions.
A.1.2 Sprague-Grundy Theorem
The Sprague-Grundy theorem provides a fundamental method for analyzing impartial games by decomposing complex games into simpler, independent subgames. As discussed in Section A.1.1, every impartial game can be represented as a directed acyclic graph (DAG). However, directly computing Grundy numbers recursively from terminal states is often impractical.
@@ A.1.2 Sprague-Grundy Theorem
Let $\mathcal{G}_{1}=(X_{1},F_{1}),\mathcal{G}_{2}=(X_{2},F_{2}),\dots,\mathcal{G}_{n}=(X_{n},F_{n})$ be DAGs representing $n$ impartial games. The disjunctive sum of $\mathcal{G}_{1},\dots,\mathcal{G}_{n}$ is a DAG $\mathcal{G}=(X,F)$ defined as follows:
@@
This result implies that Grundy numbers for complex games, such as Kayles or Chomp, can be efficiently computed by decomposing them into simpler subgames and combining the Grundy numbers using bitwise XOR. For example, consider a variant of the game Kayles played on two separate rows of pins, each forming an independent subgame. Suppose we computed the Grundy numbers separately for these rows, obtaining Grundy numbers $7$ for the first row and $4$ for the second row. By the Sprague-Grundy theorem, the combined Grundy number of the position is given by:
@@
A.1.3 Basic Discussions on the Optimal Strategy
@@
$G(S)=0\;\;\Longleftrightarrow\;\;0\notin\{G(S^{\prime})\mid S^{\prime}\in\mathcal{R}(S)\}$
where $\mathcal{R}(S)$ is the set of all states reachable from $S$. Concretely, if $G(S)=0$, then no valid move leads to a successor $S^{\prime}$ with $G(S^{\prime})=0$. In other words, from $S$, the player to move cannot transition the game into a $G(\cdot)=0$ state. Because a state $G(S^{\prime})=0$ corresponds to a losing position for the player who faces it, the mover in state $S$ has no way to force the opponent into a losing position on the next turn. Hence, $S$ is losing for the player to move.
@@
Game Rule Mathematical Strategy
NIM $k$ heaps; remove from exactly one heap At least 1 up to a fixed maximum Last move wins (normal play) Core: Nim-sum $n_{1}\oplus\cdots\oplus n_{k}$ Winning: Nim-sum $\neq 0$ (move to 0) Losing: Nim-sum =0=0 Single-heap case $\Rightarrow$ modular arithmetic simplification
Fibonacci Nim Start with $n$ items First move: $1\leq k<n$ Next: $1\leq x\leq 2\times$ (previous move) Last move wins Core: Governed by Fibonacci numbers and Zeckendorf’s theorem Zeckendorf: $m=F_{k_{1}}+F_{k_{2}}+\dots+F_{k_{r}}$, $|k_{i}-k_{j}|\geq 2$ Winning: Reduce to nearest smaller Fibonacci $F_{j}$ Losing: Pile size $F_{n}$ Optimal first move: remove the smallest Fibonacci in decomposition
Kayles Row of $n$ pins Knock down 1 pin or 2 adjacent pins Last move wins Core: Grundy numbers via Sprague–Grundy theorem Recurrence: $G(n)=\mathrm{mex}\{G(n-1),G(n-2),G(a)\oplus G(b)\}$, a+b=nka+b=n-k, $k\in\{1,2\}$ Winning: Grundy $\neq 0$ Losing: Grundy =0=0; periodic mod 12 when $n\geq 70$ Strategic principles: split into XOR-sum =0=0, mirror symmetric moves, avoid isolated single pins
Chomp $m\times n$ chocolate grid, (1,1) poisoned Choose square; remove all above/right Last move wins Core: Strategy-stealing $\Rightarrow$ first player always wins Winning: Guaranteed for first player (though constructive proof unknown) Losing: Not characterized except for small grids Analysis via posets under component-wise order Strategic principles: control antidiagonal, enforce symmetry, reduce to subgames, avoid isolated columns
Corner Queen Queen at $(x,y)$ on $m\times n$ grid Move: left, down, or diagonal left-down by $k>0$ Win by reaching $(0,0)$ Core: Equivalent to Wythoff’s Game; Beatty sequence structure Losing ($P$-positions): $(\lfloor k\phi\rfloor,\lfloor k\phi^{2}\rfloor)$, $\phi=\tfrac{1+\sqrt{5}}{2}$ Examples: $(1,2),(3,5),(4,7),(6,10),(8,13),(9,15),\dots$ Winning: Move to nearest $P$-position
@@
Table 4: MetaNIM Arena: Detailed Rules and Mathematical Strategies of Game Variants
A.2 Dataset and Simulator
We construct a simple dataset and a simulator based on the MetaNIM Arena. The dataset focuses on specific scenes in each game and does not require an opponent model, while the simulator enables interaction with an opponent that receives prompts and outputs actions, where we employ the GPT-4o model as the opponent.
@@ A.2 Dataset and Simulator
Action space 1–3 dynamic (max 30) $x,y$ coordinate single pin or
(scenario-dependent) two adjacent pins
@@
A.2.1 Game Prompts
We provide how our MetaNIM Arena game’s (NIM, Fibonacci, Chomp, Kayles, Corner-Queen) prompts are constructed as shown in Table 9. Any users can modify the games as they want to make.
Game Game Prompt
Standard #Game Role: You are {agent[‘name’]}, a participant in a game of Nim variants. #Objective: Your goal is to win the game by taking all remaining items on your turn, leaving no items for your opponent. The person who takes the last item wins. #Game Rule: There is a single pile of items. You can take between 1 and {max_take} items on your turn. #Current State: There are {remaining_items} items remaining in the pile. #Task: Based on the current state of the game, decide how many items you will take (between 1 and {max_take}) on this turn. Output Format: The output should be a Markdown code snippet with the following scheme, including leading and trailing triple backticks with "json" and: ‘‘‘{action: integer // This is an action you take. Only integer between 1 and 3.} ’’’
ReAct Output Format: The output should be a Markdown code snippet with the following scheme, including leading and trailing triple backticks with "json" and: ‘‘‘{reasoning: string // This is the reason for the action action: integer // This is an action you take based on the reasoning. Only integer between 1 and 3.} ’’’
CoT Output Format: The output should be a Markdown code snippet with the following scheme, including leading and trailing triple backticks with "json" and: ‘‘‘ { reasoning: string // This is the reason for the action action: integer // This is an action you take based on the reasoning. Only integer between 1 and 3. } ’’’ Let’s think step-by-step. What is the best move for you?
DReaMAD #Game Situation Reinterpretation: game_prompt : Below is a game description. Extract the key information. Game Description: {current state} ### Format response as: ‘‘‘{ game definition: string // What is the definition of this game? winning condition: string // How to win the game. move constraints: string // What actions are allowed per turn. } ’’’ #General Strategy Formulation: strategy_prompt : Based on the game information below, derive the general winning strategy in this game Game: {game definition} Winning Condition: {winning condition} Move Constraints: {move constraints} Current State: {current state in very short} ### Format response as: ‘‘‘{ state evaluation: string // How to assess the game state. winning strategy: string // Winning strategy in this turn to win this game. endgame tactics: string // Best strategy in a near-win situation. }’’’ #Perspective Diversification: Refine the initial game prompt to improve decision-making based on the Game and Strategy Information. Initial Prompt: {given initial prompt} Game and Strategy Information: Game: {game definition} Strategy: - State Evaluation: {state evaluation} - Winning Strategy: {winning strategy} - Endgame Tactics: {endgame tactics} ### Format response as: ‘‘‘{ optimized prompt: string // The refined prompt that clearly directs decision-making.} ’’’
Table 8: Standard, ReAct, CoT, DReaMAD Prompt in NIM game.
Game Game Prompt
NIM #Game Role: You are {agent[’name’]}, a participant in a game of Nim variants. #Objective: Your goal is to win the game by taking all remaining items on your turn, leaving no items for your opponent. The person who takes the last item wins. #Game Rule: There is a single pile of items. You can take between 1 and {max_take} items on your turn. #Current State: There are {remaining_items} items remaining in the pile. #Task: Based on the current state of the game, decide how many items you will take (between 1 and {max_take}) on this turn.
Fibonacci #Game Role: You are {agent[’name’]}, a participant in a simple Fibonacci game. #Objective: Your goal is to win the game by taking all remaining stones on your turn, leaving no stones for your opponent. The person who takes the last stones wins. #Game Rule: 1. There is a single pile of stones.
@@ A.2.1 Game Prompts
#Current State: There are {remaining_items} stones remaining in the pile. #Task: You are the first player. Based on the current state of the game, decide how many items you will take (between 1 and {remaining_items - 1}) on this turn.
Chomp #Game Role: You are {agent[’name’]}, a participant in a game of Chomp. #Objective: Your goal is to force your opponent to take the top-left corner of the grid (position (0, 0)). #Game Rule: 1. The game is played on a square grid.
2. On your turn, you select a position (row, col).
4. The player forced to select (0, 0) loses.
#Current State: The grid is represented as a binary matrix, where ’1’ means the position is still available, and ’0’ means it is removed: {remaining_grid} #Task: Based on the current state of the grid, decide which position (row, col) you will select.
Kayles #Game Role: You are {agent[’name’]}, a participant in a game of Kayles. #Objective: Your goal is to win the game by leaving your opponent with no valid moves. The player who takes the last pin(s) wins. #Game Rule: 1. There is a single row of pins.
2. On your turn, you can remove:
• 1 pin,
• 2 adjacent pins.
3. You cannot remove non-adjacent pins or pins that have already been removed.
#Current State: The row of pins is represented as a binary string: